EVA Report – May 22nd

Crew 195 EVA report 22.05.18


EVA Crew Members –

Participants: Scott D., Mason, Christian, Rae, Ruslana, Dana Levin, Barry

Participants remaining in Hab: Chris Davis and Madison Ridgeway


Destination: Gateway to Lith

EVA Objectives

1. Simulation of medical contingency scenarios during EVA.


Successful contingency training for medical sim



ATVs/Rovers Used

ATVs x 3, EVs x3 (Spirit, Opportunity, Deimos)

Journalist Report- May 22nd

[tittle Journalist Report – May 22]

MDRS Crew 195
Sol 2

Dubbed Nerd Day, our EVA centred around finding evidence of life on Mars using the drone. It was a longer ride on the quads and buggies (gee life is tough) and a few short strolls up and down the gentle hills en route.

Even with the wind picking up and a spattering of Martian rain, the drone buzzed about, up and down the wash and sticky-beaking on the crew. Hey "Lead", do you ever get that feeling you’re being watched?

We’ve said goodbye to one crew member today and will gain more on the resup ship. Afternoon downtime was filled with a bit of reading, electronic Chess games, snoozing, exercise in the Engineering tunnel, trading "war stories" and rehydrating freeze dried blueberries with orange Gatorade. As you do.

For tonight’s entertainment: The Martian. Most of us have seen it before but this time it will be with new eyes. Freshly ultrasound gelled eyes.

Sol Summary – May 22nd

[tittle Sol 2 Summary – May 22]

Crew 195 Daily Summary Report 22.05.2018

MDRS Daily Summary Report for sol 2

Summary Title

EVA Bliss.

Mission Status:


Sol Activity Summary:

Second EVA to Gateway of Lith. Successful sim of medical resuscitation

Look Ahead Plan: We will return to Gateway to Lith to map area.

Anomalies in work:

ATV 300- continued issues with battery. Unable to hold charge.

Weather:Cloud cover for most of day. Pleasantly cool.

Crew Physical Status:Doing well.

EVA: Tomorrow

Reports to be file:

Operations Report

Journalist report

Photo of the day

Support Requested:


Operations Report – May 22nd

[tittle Operations Report – May 22]

Crew 195 Operations Report 22.05.18

SOL: 2

Name of person filing report: Chris Davis

Non-nominal systems: Continuing run Diesel generator at night. No service rep for Solar yet. Septic serviced today.

Notes on non-nominal systems:

Generator (hours run): Overnight approximately 10 plus 4 hours today per Shannon

Solar— SOC 0%

Diesel – Over 50%

Propane – 51 PSI

ATV Fuel (Ethanol free gasoline) – 4 gallons.

Water (trailer) – 500 gallons.

Water (static) – Approximately 150

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 1351176

Toilet tank emptied: Yes
Date last emptied: 5.22.18

ATV’s (Honda, 300, 350.1, 350.2, 350.3): all except Honda and 300

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: Approximately 1 gallons

# Hours the ATVs were Used today: 1 hours

Notes on ATVs: Honda and ATV 300 out of service.

Deimos rover used: yes
Hours: N/A
Beginning charge: 100%
Ending charge: 74 %
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: Yes

Beginning charge: 100 %
Ending charge: 76%
Currently charging: Yes

Opportunity rover used: No
Beginning charge: NA
Ending charge: NA
Currently charging: NA

Curiosity rover used: No
Beginning charge: 100 %
Ending charge: 71%
Currently charging: Yes

Hab Car used?: No
Where was Hab Car used?:
Why was Hab Car used?:

General notes and comments:

Summary of internet: Working well

Summary of suits and radios: working well

Summary of Hab operations: No problems

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of health and safety issues: none noted

Questions, concerns and requests to Mission Support: None.

Copyright © The Mars Society. All rights reserved. | Main Site